OpenStudy (anonymous):

If the input voltage from the transformer is 32 V, what is the output voltage? (output has 60 turns of wire, input has 4 turns of wire.)

OpenStudy (anonymous):

i know its not math, but i cant figure it out and i cant find the answer anywhere else.

OpenStudy (anonymous):

5 mins and you will get your answer :)

OpenStudy (anonymous):

every transformer contains primary winding and a secondary winding , here Vp=32 V and (voltage of the primary winding ) Np =4 (turns of primary winding) Vs= ? (wer trying to find it voltage of the secondary) , Ns =60 ( turns of the secondary) we have this equations Vp*Np=Vs*Ns we have Vp,Np,Nv //trying to find Vs 32*4=60Vs Vs = ............

OpenStudy (anonymous):

32*4 = 128 so i divided 128 by 60 and got 2.13. so would that be the answer?